Write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b
(4,7);5x+y=9

Answer by rothauserc(4718)

5x + y = 9
:
y = -5x +9
:
the slope of this line is -5, therefore the perpendicular line is the negative reciprocal of -5 which is 1/5
:
y = x/5 +b
:
to find b, we use the point (4,7)
:
7 = 4/5 +b
:
b = 7 -4/5 = 35/5 -4/5 = 31/5
:
*****************************
y = x/5 +31/5
*****************************
: